### v3

# Semi-supervised 3D Semantic Scene Completion with 2D Vision Foundation Model Guidance

Abstract
Accurate prediction of 3D semantic occupancy from 2D visual images is vital in enabling autonomous agents to comprehend their surroundings for planning and navigation. State-of-the-art methods typically employ fully supervised approaches, necessitating a huge labeled dataset acquired through expensive LiDAR sensors and meticulous voxel-wise labeling by human annotators. The resource-intensive nature of this annotating process significantly hampers the application and scalability of these methods. We introduce a novel semi-supervised framework to alleviate the dependency on densely annotated data. Our approach leverages 2D foundation models to generate essential 3D scene geometric and semantic cues, facilitating a more efficient training process. Our framework exhibits notable properties: (1) Generalizability, applicable to various 3D semantic scene completion approaches, including 2D-3D lifting and 3D-2D transformer methods. (2) Effectiveness, as demonstrated through experiments on SemanticKITTI and NYUv2, wherein our method achieves up to 85% of the fully-supervised performance using only 10% labeled data. This approach not only reduces the cost and labor associated with data annotation but also demonstrates the potential for broader adoption in camera-based systems for 3D semantic occupancy prediction.
Introduction
Camera-based 3D semantic scene completion (SSC) is an important vision task with many applications in both indoor and outdoor scenarios (Song et al. 2017; Liu et al. 2018; Roldao, de Charette, and Verroust-Blondet 2020; Wu et al. 2020; Li et al. 2020; Dourado et al. 2021; Huang et al. 2023b; Cao and de Charette 2022; Tong et al. 2023; Yao et al. 2023; Li et al. 2023; Zhang, Zhu, and Du 2023). Given an input image, we aim to predict a complete 3D voxel representation of volumetric occupancy and semantic labels for a scene. This task is challenging due to the lack of 3D information from 2D images. Most SSC methods (Song et al. 2017; Li et al. 2023; Zhang, Zhu, and Du 2023) solved this by learning in a fully supervised manner, i.e., relying on dense 3D SSC annotations. However, obtaining labels requires setting up a vehicle with LiDAR sensors and meticulous human annotations for each voxel, limiting the scalability and deployment of many SSC systems. An alternative approach to mitigate the annotation cost associated with SSC involves leveraging self-supervised learning (Cao and de Charette 2023; Huang et al. 2023a; Zhang et al. 2023a) which only requires unlabeled images for training. Nevertheless, these methods output poor results with a large performance gap compared to fully supervised counterparts.
Figure 1: Visual comparison on SemanticKITTI: VFG-SSC surparses self-supervised methods like SelfOcc (Huang et al. 2023a) by 4% mIoU and outperforms other semi-supervised methods: Self-Training (Lee et al. 2013) and Mean Teacher (Tarvainen and Valpola 2017) by 1.5% mIoU.
To address these limitations, this paper introduces the problem of semi-supervised camera-based 3D semantic scene completion (Semi-SSC). In this setup, a small percentage of images are labeled with 3D semantic occupancy, while the majority remain unlabeled. To overcome the challenge of limited labeled data, semi-supervised learning approaches like Self-Training (Lee et al. 2013) and Mean-Teacher (Tarvainen and Valpola 2017) can be employed. In Self-Training (ST), a fully supervised SSC network first learns from labeled data and then predicts pseudo-labels for unlabeled data. These pseudo-labels are used to retrain the network, leveraging the unlabeled data. In contrast, the Mean-Teacher (MT) approach involves a more stable Teacher network, derived from the student network’s Exponential Moving Average (EMA), which predicts pseudo-labels for unlabeled images. These pseudo-labels are combined with labeled data to continually train the student network. While there are advanced semi-supervised learning techniques for 3D, such as LaserMix (Kong et al. 2023), CutMix (Yun et al. 2019), and ClassMix (Olsson et al. 2021), these methods rely on a one-to-one correspondence between input and output or homogeneous data. This approach does not apply to Semi-SSC, where the input consists of 2D images and the output is in 3D voxels.
On the other hand, recent vision foundation models (VFMs) have shown the ability to robustly extract metric depth (Piccinelli et al. 2024) and semantics (Zhang et al. 2023b) from unlabeled images. However, these capabilities are not yet leveraged in SSC. To bridge this gap, we introduce VFG-SSC, a method that leverages 2D VFMs to generate 3D priors from unlabeled images, thereby improving the quality of pseudo-labels during training. Adapting 2D VFM knowledge to 3D presents challenges such as depth scale ambiguity and feature misalignment from segmentation models. To address these, we developed a novel attention-based enhancement module that (1) globally refines features and (2) maintains linear complexity, enabling efficient processing of large 3D datasets. Futhermore, we propose a technique to accumulate VFM information over image sequences, reducing noise and biases compared to using a single image. Our method can serve as a universal plugin for any camera-based SSC method, enabling a semi-supervised alternative for fully supervised approaches.
We evaluate our approach on the outdoor SemanticKITTI (Behley et al. 2019) and indoor NYUv2 (Silberman et al. 2012) datasets. Our results demonstrate strong performance across both settings, achieving 85% of the accuracy of fully supervised methods while utilizing only 10% of the 3D semantic occupancy annotations.
In summary, our contributions are as follows:
  • We introduce a new semi-supervised 3D semantic scene completion (Semi-SSC) framework to reduce the high cost of 3D voxel-wise annotation.
  • We propose VFG-SSC, a novel approach that leverages 3D cues from 2D vision foundation models for depth and segmentation.
  • We design a new enhancement module to effectively combine 3D cues with 3D features, seamlessly integrating with existing SSC methods.
  • Our pipeline achieves 85% of fully supervised performance using just 10% of the labeled data.
Related Work
This section reviews literature on fully, semi, and self-supervised 3D semantic scene completion, followed by an overview of the 2D depth and semantic segmentation foundation models used in our framework.
3D Semantic Scene Completion (SSC). The primary goal of 3D semantic scene completion (SSC) or 3D semantic occupancy prediction is to jointly infer the geometry and semantics of a scene. Initially introduced in SSCNet (Song et al. 2017) for indoor scenes like NYUv2 (Silberman et al. 2012), the task has since been extended to outdoor environments such as SemanticKITTI (Behley et al. 2019) (Roldao, de Charette, and Verroust-Blondet 2020; Li et al. 2020; Yan et al. 2021; Cheng et al. 2021; Mei et al. 2023; Xia et al. 2023; Jia et al. 2023). Camera-based SSC methods, which are cost-effective and easily deployable, have gained traction. MonoScene (Cao and de Charette 2022) was the first to use only RGB images for SSC by projecting 3D voxels onto 2D images and processing them with a 3D UNet to predict 3D semantic occupancy. VoxFormer (Li et al. 2023) improves upon this by leveraging depth estimation priors to focus cross-attention on voxels near object surfaces. OccFormer (Zhang, Zhu, and Du 2023) further enhances 3D feature encoding using windowed attention on local and global transformer pathways and employs a mask decoder, similar to Mask2Former (Cheng et al. 2022), to predict 3D SSC. However, these methods still rely on expensive voxel-wise annotations, limiting their practical application.
Self-supervised SSC. One way to reduce the annotation cost of SSC is through self-supervised learning, which leverages video sequences along with predicted depth and semantic segmentation as pseudo-labels. Methods like SceneRF (Cao and de Charette 2023), SelfOcc (Huang et al. 2023a), and OccNeRF (Zhang et al. 2023a) use volumetric rendering on predicted 3D semantic occupancy to generate RGB images, depth maps, and segmentation maps, allowing for loss computation with pseudo-labels. Although self-supervised SSC requires less supervision than our VFG-SSC approach, its performance is significantly lower than that of semi-supervised and fully-supervised methods, making it unsuitable for practical applications.
Semi-supervised learning for 2D and 3D LiDAR segmentation. Semi-supervised image segmentation methods (Chen et al. 2021; Liu et al. 2022; Kwon and Kwak 2022; Wang et al. 2022; Yang et al. 2023; Hoyer et al. 2023) have achieved strong performance on autonomous datasets using two key strategies: consistency regularization and entropy minimization. Consistency regularization applies augmentations like CutMix (Yun et al. 2019) and ClassMix (Olsson et al. 2021) to ensure stable predictions with perturbed inputs, while entropy minimization reduces uncertainty in classifications for unlabeled data, leading to more confident predictions. In 3D LiDAR segmentation, LaserMix (Kong et al. 2023) uses mixing schemes for consistency enforcement, and methods like 3DIoUMatch (Wang et al. 2021) and DetMatch (Park et al. 2022) use multi-sensor data to generate high-quality pseudo-labels for object detection. These methods require homogeneous input-output, Semi-SSC is more challenging as it takes a 2D image as input but outputs 3D semantic occupancy. Therefore it is non-trivial to apply the above methods to our task. Furthermore, Semi-SSC only leverages images for unlabeled data, as using multi-sensors can infer significant setup costs for data collection.
2D foundation models for depth estimation and segmentation. Segmentation models like Segment Anything (Kirillov et al. 2023) have introduced class-agnostic segmentation, demonstrating robustness across diverse datasets. Following this, models such as X-Decoder (Zou et al. 2023), SEEM (Zou et al. 2024), and OpenSEED (Zhang et al. 2023b) have emerged as general-purpose panoptic segmentation tools, offering zero-shot application on autonomous datasets. For depth estimation, models like Metric3Dv2 (Yin et al. 2023) and UniDepth (Piccinelli et al. 2024) excel in robust metric depth estimation by leveraging large-scale datasets. In our work, we utilize these 2D VFMs to generate 3D clues, addressing the challenge of limited labeled data.
Figure 2: The overall architecture of the proposed VFG-SSC network. The model generates 3D clues from the 2D foundation models to enhance the inferred 3D features volume and produces the final 3D semantic occupancy grid. We train a model using all frames (solid red box) to generate pseudo-labels of the unlabeled data.
Semi-Supervised Semantic Scene Completion
Problem Setting
Given a sequence $\mathbf{Q}_{t}=\{I_{t-k},I_{t-k+1}\ldots,I_{t-1},I_{t}\}$ of $k$ consecutive frames, the SSC model $f_{\theta}$ generates a semantic occupancy grid which is defined in the coordinate system of the ego vehicle at the timestamp $t$. Each voxel of the grid is categorized as either empty or occupied by a specific semantic class. The grid can be obtained as follows: $\mathbf{\hat{Y}}_{t}=f_{\theta}(\mathbf{Q_{t}})$ where $\mathbf{\hat{Y}}_{t}\in\mathbb{R}^{H\times W\times Z\times(C+1)}$. $H$, $W$, and $Z$ denote the voxel grid’s height, width, and depth, and $C$ is the number of the semantic classes. In a semi-supervised setting (Semi-SSC), the dataset contains two non-overlapping subsets:
  • Labeled data $\mathcal{D}^{L}=\{(\mathbf{Q}_{t},\mathbf{Y}_{t})\}_{t=1}^{L}$ where each image $I_{t}$ has a corresponding 3D occupancy ground-truth $\mathbf{Y}_{t}$.
  • Unlabeled data $\mathcal{D}^{U}=\{\mathbf{Q}_{t}\}_{t=1}^{U}$, which only contains images with no LiDAR and the amount of the unlabeled data is significantly larger than labeled data, i.e, $U\gg L$.
Methods Precision $\uparrow$ Recall $\uparrow$ mIoU $\uparrow$
Sup-only 47.53 47.09 9.16
Self-Training (ST) 42.83 59.35 9.48
3D clues as pseudo-labels 48.60 26.19 9.27
Our approach 48.01 58.36 11.73
Table 1: Comparison of Semi-SSC approaches on 5% labeled data on the SemanticKITTI val set. The Sup-only baseline does not use pseudo labels from unlabeled data during training.
Our Customized Self-Training for Semi-SSC
A common strategy for tackling the semi-supervised problem is Self-Training. This involves first training a supervised model $f_{\theta}$ on labeled data $\mathcal{D}^{L}$ and then generating pseudo-labels for the unlabeled dataset $\mathcal{D}^{U}$. After that, the model $f_{\theta}$ is retrained using both the labeled and pseudo-labeled data. Based on this Self-Training approach, our key contribution is to enhance the quality of the pseudo-label by incorporating 3D priors extracted from 2D vision foundation models. We now outline our three-step process in detail.
Figure 3: Qualitative of Semi-SSC approaches. With Sup-only and ST, the scene layout is reasonably reconstructed, however, dynamic object prediction is incorrect due to many false positives. With 3D clues as pseudo-labels, the cars are correctly predicted but occluded regions do not have a prediction (assigned as empty class). Our methods obtain reliable object predictions (high precision) and reasonable reconstruction in occluded space (high recall).
  • Step 1 - Training on labeled data: We start by extracting efficient 3D cues from 2D images using 2D foundation models. These cues are then combined with the 3D features extracted from a 3D SSC network $f_{\theta}$ through our proposed enhancement module $g_{\phi}$. The networks $f_{\theta}$ and $g_{\phi}$ are jointly trained on labeled data $\mathcal{D}^{L}$, enabling the proposed enhancement module to learn the efficient fusion between 3D clues and 3D features with the supervision of labeled data.
  • Step 2 - Generating pseudo-label for unlabeled data: The trained $f_{\theta}$ and $g_{\phi}$ from Step 1 are used to make prediction on unlabeled data $\mathcal{D}^{U}$, resulting in a pseudo-labeled dataset $\hat{\mathcal{D}}^{L}=\{(\mathbf{I}_{t},\hat{\mathbf{Y}}_{t})\}_{t=1}^{U}$.
  • Step 3 - Retraining with labeled and pseudo-labeled data: Finally, we retrain the original SSC model $f_{\theta}$ using the combined dataset $\mathcal{D}^{L}\cup\hat{\mathcal{D}}^{L}$ with supervised segmentation losses. Note that only $f_{\theta}$ is utilized during test-time; 3D cues and $g_{\phi}$ are not used at this stage.
In the next section, we focus on presenting our main contributions in Step 1 and 2.
Our VFG-SSC Network
To highlight the importance of improving pseudo-label quality in Self-Training, we conducted a quick test by training $f_{\theta}$ with various methods shown in Tab. 1. As can be seen, using pseudo labels achieves lower precision compared to models trained on 5% labeled data, primarily boosting recall. While the model can infer plausible geometry in occluded regions, relying solely on image input introduces ambiguity. To overcome this limitation, we incorporate 3D cues from 2D VFMs as pseudo labels for the unlabeled data. These 3D cues offer higher precision but lower recall, as they provide accurate yet sparse estimates of depth and semantics. By combining these sources, we enhance both precision and recall, leading to an overall improvement in mIoU. This approach motivates our strategy to align 3D cues with Self-Training by integrating them as inputs during pseudo-label generation. We achieve this enhancement through an attention-based mechanism that effectively fuses the two feature sets.
Our VFG-SSC network, as shown in  Fig. 2, includes a unique 3D clue generation module, a semantic scene completion network $f_{\theta}$, and an enhancement module $g_{\phi}$.
3D Clues Generation
As mentioned in the previous section, we exploit 2D VFMs of depth estimation and semantic segmentation to create 3D clues. A 3D clue is defined as a discrete set of voxels $\mathbf{Y}_{\text{clues}}\in\mathbb{R}^{H\times W\times Z\times(C+1)}$. A straightforward approach is to extract depth and semantic information from the current image, and then use the depth data to project the segmentation into a 3D voxel representation. However, as shown in  Fig. 4, 3D clues generated from a single frame often contain numerous irrelevant artifacts that do not align with the actual scene. To address this issue, we propose generating clues using temporal image sequences. Particularly, given a set of $N$ input images of a scene, we extract depth maps and semantic segmentation map from VFMs. For each pixel $u=[u_{x},u_{y}]$ and its corresponding depth $d$, we un-project the semantic map $u$ to 3D point $x$ in world coordinate by computing $x=E^{-1}\times K^{-1}\times[u_{x},u_{y},d]^{T}$, where $K$ and $E$ are the camera intrinsic and extrinsic matrices respectively. This results in a set of semantic point clouds $X\in\mathbb{R}^{N\times 4}$ (3D coordinates and semantic class). We convert these point clouds to voxels to obtain $\mathbf{Y}_{\text{clues}}$.
Figure 4: 3D clues visual comparisons on different settings: using only the current image and accumulated 3D clues from temporal frames with and without filtering.
As shown in  Fig. 4, accumulating 3D clues from temporal frames (Multi Frame) results in a more accurate scene representation compared to a single frame, however, dynamic classes like vehicles become indistinguishable. To tackle this, we introduce a regularization pipeline that leverages semantic and geometric constraints. We begin by segmenting the semantic point cloud produced in the previous step by class and applying heuristic filters: a radius filter to remove outliers based on the density of nearby points, and a statistical filter to eliminate points that deviate significantly from the average distance. Results in Tab. 8 show that this pipeline is crucial to effectively utilize VFMs. A pseudo-code implementation is provided in the Supplementary Material.
3D Clue-Guided Enhancement Module
This module, $g_{\phi}$, fuses the 3D clues from the previous module with the 3D features generated by the scene completion network $f_{\theta}$ to create an enhanced 3D feature volume. A naive approach is to concatenate these 3D clues with the features from $f_{\theta}$ to obtain a refined feature. However, our visualized results in Fig. 3, reveal that these features contain complementary information that can be better utilized. The 3D clues can reliably estimate the semantics and geometry of visible surfaces, but due to limited camera FoV, many voxels remain without sufficient information. Concatenation alone may introduce ambiguities in these regions. An alternative approach involves cross-attention, which associates each feature in $f_{\theta}$ with all 3D clues, but this can lead to a significant increase in computational cost. To balance effectiveness and efficiency, we propose using a local attention mechanism to merge the two features, achieving both global context aggregation and linear computational complexity.
Our enhancement module uses a set of self-attention blocks to align two features together implicitly. Specifically, given 3D clues $\mathbf{Y}_{\text{clues}}$ and 3D features $\mathbf{Y}_{\text{feats}}$, we project discretized 3D clues into continuous embeddings. We then concatenate the aligned features with the original 3D features and use self-attention for enhancement. Note that, we adopt Dilated Neighborhood Attention (Hassani et al. 2023), a recently proposed method with linear time and space complexity to self-attention to reduce the complexity.
Backbone Method 1% 5% 10%
[…]
SceneRF Self-supervised N/A mIoU, 13.84 IoU
SelfOcc Self-supervised 5.02 mIoU, 21.97 IoU
SelfOcc Finetuned 5.55 29.57 7.5 32.43 8.3 33.66
OccFormer Fully-Supervised 14.08 mIoU, 36.04 IoU
Supervised-only 6.59 26.31 9.16 32.73 10.45 34.25
Mean-Teacher 6.76 26.54 9.41 31.67 10.53 33.28
Self-Training 6.75 26.61 9.48 33.12 10.49 33.94
VFG-SSC (Ours) 9.40 30.40 11.73 35.76 12.38 35.25
VoxFormer Fully-Supervised 14.40 mIoU, 43.50 IoU
Supervised-only 6.35 34.04 9.56 39.68 10.52 40.28
Mean-Teacher 7.20 38.97 9.58 37.70 10.56 38.66
Self-Training 7.60 35.40 10.51 39.83 11.06 41.05
VFG-SSC (Ours) 9.32 36.78 11.15 39.96 12.19 41.57
Table 2: Quantiative results on the validation set of the SemanticKITTI (Behley et al. 2019) dataset. Faded rows denote results for reference purposes only, not direct comparison.
Experiments
Datasets and metrics: We utilize two 3D semantic occupancy datasets, SemanticKITTI (Behley et al. 2019) and NYUv2 (Silberman et al. 2012), which encompass outdoor and indoor environments, respectively, to establish new semi-supervised benchmarks. For SemanticKITTI, we sample 40, 198, and 383 frames (corresponding to 1%, 5%, and 10% of the training set), consistent with existing setups (Wang et al. 2023; Behley et al. 2019). For NYUv2, we uniformly sample 40 and 80 frames (representing 5% and 10% of the training set). We evaluate performance using mIoU for semantic class and IoU for binary occupancy predictions.
Implementation details: For SemanticKITTI, we use OccFormer (Zhang, Zhu, and Du 2023) and VoxFormer (Li et al. 2023) as our primary SSC networks ($f_{\theta}$). VoxFormer employs a two-stage training pipeline: we pre-train the query proposal with 1/5/10% labeled data in stage 1, then apply VFG-SSC in stage 2. For OccFormer, which uses additional depth supervision from LiDAR, we supplement the missing LiDAR data in unlabeled sets with pseudo-LiDAR generated from a depth model (Piccinelli et al. 2024). We maintain consistent settings across both models for fair comparison. For NYUv2, we use MonoScene (Cao and de Charette 2022) and OccDepth (Miao et al. 2023) as the SSC networks.
UniDepth (Piccinelli et al. 2024) is used for depth estimation, and SEEM (Zou et al. 2024) for semantics, serving as our foundational models. We employ 2 layers of Dilated Neighborhood Attention with 4 heads, a kernel size of 7, and 4 dilation rates (1, 2, 4, 8). Additional details on losses for training each SSC network and further implementation specifics are provided in the Supplementary Material.
[…]
road
sidewalk
parking
other-grnd
building
car
truck
bicycle
motorcycle
other-vehi
vegetation
trunk
terrain
person
bicyclist
motorcyclist
fence
pole
traffic-sign
MonoScene 11.1 34.2 54.7 27.1 24.8 5.7 14.4 18.8 3.3 0.5 0.7 4.4 14.9 2.4 19.5 1.0 1.4 0.4 11.1 3.3 2.1
OccFormer 13.2 33.7 57.3 31.1 29.7 13.7 16.8 21.5 4.5 2.4 2.2 3.2 15.5 4.3 22.5 2.7 2.3 0.0 12.5 4.1 4.2
VoxFormer 14.3 42.0 56.5 30.0 25.8 12.3 23.7 23.4 5.1 3.4 1.6 2.7 24.5 8.6 24.4 1.6 1.3 0.0 15.4 6.3 6.0
VFG-SSC (Occ) 11.1 30.5 55.2 26.7 25.6 0.0 14.8 20.5 1.6 2.3 2.2 5.8 13.1 3.6 20.7 3.4 0.0 0.0 10.5 3.4 2.1
VFG-SSC (Vox) 11.8 40.6 55.5 23.8 16.3 0.0 20.9 21.5 2.2 1.4 1.8 3.5 23.7 8.2 20.0 1.5 0.0 0.0 12.5 5.8 4.9
Table 3: Quantitative comparisons between our method and other fully-supervised methods on the hidden testing set of SemanticKITTI (Behley et al. 2019) dataset. Faded rows denote results for reference purposes only, not direct comparison.
Comparison with Related Methods
As the first work to address the 3D semantic scene completion task in a semi-supervised manner, we adapt two semi-supervised learning approaches as baselines: Mean Teacher (Tarvainen and Valpola 2017) , which leverages consistency regularization, and Self-Training (Lee et al. 2013) with entropy minimization. In the Semi-SSC setting, we train our model using only 1%, 5%, and 10% of the labeled training data and compare its performance against other methods trained with the same data percentages. For reference, we also include results from fully supervised training using 100% labeled data, as well as two self-supervised methods, SceneRF (Cao and de Charette 2023) and SelfOcc (Huang et al. 2023a), which rely entirely on unlabeled data. Additionally, we evaluate a version of SelfOcc fine-tuned with the same labeled data percentages by incorporating a segmentation head.
SSC Network Method 5% 10%
[…]
MonoScene Fully-Supervised 26.94 mIoU, 42.51 IoU
Supervised-only 13.36 29.17 17.17 34.68
Self-Training 14.00 30.64 17.32 35.20
VFG-SSC 18.23 43.35 22.20 44.37
OccDepth Fully-Supervised 29.03 mIoU, 44.17 IoU
Supervised-only 16.69 30.97 19.01 37.83
Self-Training 16.72 30.49 19.42 38.07
VFG-SSC 22.54 43.61 23.92 44.85
Table 4: Quantiative results on the testing set of the NYUv2 (Silberman et al. 2012) dataset. Faded rows denote results for reference purposes only, not direct comparison.
Input Mean-Teacher Self-Training SelfOcc Ours Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 5: Qualitative results on the validation set of the SemanticKITTI (Behley et al. 2019) dataset.
Input Sup-only Self Training Ours Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 6: Qualitative results on the testing set of the NYUv2 (Silberman et al. 2012) dataset.
Quantitative results: For SemanticKITTI, we present IoU and mIoU in  Tab. 2. Additionally, we validate our results by reporting VFG-SSC performance with a 10% labeled data setting on the SemanticKITTI hidden test set in Table  Tab. 3. Our VFG-SSC significantly outperforms baseline methods. Remarkably, our approach achieves comparable results to fully supervised counterparts, with just a 15% performance gap using 10% labeled data. It is worth noting that on the 1% setting, Mean Teacher obtains a better IoU score on VoxFormer. Therefore, a robust Semi-SSC method should achieve excellent performance in both geometric and semantic completion. Moreover, on the SemanticKITTI hidden test set, our method compares favorably with some fully supervised methods like MonoScene, despite utilizing only 10% labeled occupancy annotation. This emphasizes the effectiveness of our 3D clues and enhancement module, which is effective in generating high-quality pseudo-labels for training any SSC backbones. For NYUv2, we report the results in  Tab. 4. We consistently outperform other methods with both 5% and 10% of training data, underscoring the superiority of our approach over strong baselines. These results indicate that VFG-SSC is generalizable to different architectures, can be applied to various labeled settings, and applies to both outdoor and indoor scenarios.
Qualitative results: We present the qualitative results of our method, VFG-SSC, compared to baseline models on the SemanticKITTI and NYUv2 datasets in Fig. 5 and Fig. 6 respectively. In outdoor scenarios, VFG-SSC outperforms the baselines, particularly in accurately predicting various instance classes like cars and people. On the other hand, other baselines mis-segment persons and produce irregular car shapes (row 1). Additionally, VFG-SSC excels in capturing the broader scene layout, including long-range structures such as crossroads, sidewalks, and buildings (rows 2, 3). In indoor scenes, VFG-SSC effectively reconstructs global scene layouts, including both visible and occluded regions (row 2), while also recovering finer details of smaller objects, such as chairs, tables, and sofas (rows 1, 3).
Ablation Study
In this section, we ablate our approach on 5% labeled data of SemanticKITTI with OccFormer as the SSC network $f_{\theta}$.
Enhancement techniques: In Tab. 5, we compare our attention-based alignment method against several enhancement techniques, including simple concatenation, BEVFusion (Liang et al. 2022), and OpenOccupancy’s Adaptive Fusion (Wang et al. 2023). Additionally, we include an attention-only baseline where 3D features and 3D clues are concatenated along the feature dimension, followed by Deformable Self-Attention (Zhu et al. 2020). We evaluate the quality of the pseudo-labels generated using training ground truth (GT) labels (in step 2) and assess the final model’s predictions against validation GT labels. The results demonstrate that our method consistently outperforms the other techniques on both the training and validation splits. Notably, even with the additional clues from VFMs, simply integrating them into an SSC network does not result in significant performance improvements compared to baselines without VFMs. We attribute the substantial performance gain to our enhancement module, which effectively leverages these additional clues.
Enhancement Train Split Val Split
Methods mIoU $\uparrow$ IoU $\uparrow$ mIoU $\uparrow$ IoU $\uparrow$
Concat 13.03 42.29 10.01 34.05
BEVFusion 12.80 42.15 10.00 34.72
OpenOcc 12.42 40.08 10.37 34.37
Deformable 13.84 37.40 10.04 34.30
Ours 15.59 42.73 11.73 35.76
Table 5: Quantitative results of different fusion methods.
Depth Train Split Val Split
Model mIoU $\uparrow$ IoU $\uparrow$ mIoU $\uparrow$ IoU $\uparrow$
Depth-Anything 12.63 37.66 10.38 35.62
Metric3Dv2 14.86 41.28 11.09 36.20
UniDepth 15.59 42.73 11.73 35.76
Table 6: Study on different depth foundation models.
Study on different depth estimation models: We conducted experiments using various depth foundation models to generate 2D depth maps. Since the effectiveness of our method hinges on the accuracy of pseudo-labels, employing a more reliable depth predictor can potentially enhance the quality of the generated 3D clues. As shown in  Tab. 6, our approach achieves the best results with UniDepth. It’s worth mentioning that advancements in depth estimation will further enhance our method’s performance. In this experiment, we used SEEM (Zou et al. 2024) as the segmentation model.
Study on different segmentation models: We present a study on various semantic foundation models like SEEM (Zou et al. 2024) (pretrained on COCO), OpenSEED (Zhang et al. 2023b) (pretrained on COCO, Object365), and SegFormer (Xie et al. 2021) (pretrained on CityScapes). The comparison results reported in  Tab. 7 reveal marginal distinctions among these aforementioned models. SegFormer obtains the best results as the domain between training (CityScapes) and testing (SemanticKITTI) data is not substantial for 2D segmentation. This suggests that our method consistently produces high-quality 3D cues across diverse segmentation foundation models.
Semantic Train Split Val Split
Model mIoU $\uparrow$ IoU $\uparrow$ mIoU $\uparrow$ IoU $\uparrow$
OpenSEED 14.80 40.90 11.40 36.68
SEEM 15.59 42.73 11.73 35.76
SegFormer 16.18 41.74 11.94 36.78
Table 7: Study on different semantic foundation models.
Impact analysis of temporal and regularization: Our analysis, detailed in  Tab. 8, reveals several insights. In rows 1 and 2, we compare the performance of ST with and without the integration of a single frame of regularized 3D clues, noting a modest improvement of +0.09 mIoU. This enhancement can be attributed to the additional sparse information provided by the VFMs. In row 3, we evaluated the impact of incorporating temporal frames into 3D clue generation while omitting the regularization module, which resulted in a decrease of -0.62 mIoU. This drop indicates that the added temporal information from VFMs without regularization, introduces too much noise. However, when combining temporal accumulation with our regularization, we achieved a significant improvement of +2.25 mIoU. This demonstrates that while temporal information is valuable for VFG-SSC, regularization is essential for optimizing performance.
Train Split Val Split
Regularize Temporal mIoU $\uparrow$ IoU $\uparrow$ mIoU $\uparrow$ IoU $\uparrow$
8.37 31.6 9.48 33.12
9.28 33.87 9.57 34.26
8.61 33.16 8.86 33.64
15.59 42.73 11.73 35.76
Table 8: Impact of VFG-SSC’s temporal regularization.
Conclusion
In this paper, we have introduced a new approach for semi-supervised 3D semantic scene completion. Our method effectively utilizes 3D clues derived from 2D depth and segmentation foundation models and introduces an innovative fusion module to integrate these clues with 3D features, ensuring compatibility with most existing SSC techniques. Extensive experiments have demonstrated the effectiveness of our approach in both outdoor and indoor environments, achieving a minimal performance gap of just 15% compared to fully supervised methods while using only 10% of 3D semantic occupancy annotations. To the best of our knowledge, this is the first exploration of semi-supervised 3D SSC.
Limitations. Given the reliance on the usage of 3D clues from the 2D VFMs, our approach might not work well under challenging conditions (e.g., night-time, foggy, rainy) where the 2D VFMs fail to produce good results. Furthermore, a large domain gap between training and testing data might also harm our method, as our approach only uses a small amount of labeled data to train in the first phase.